1. Simplify the Boolean expression (A+B+C)(D+E) + (A+B+C)(D+E) and choose the best answer. 1. A+B+C 2. D+E 3. ABC 4. DE
Answer #1

PLEASE UPVOTE THE SOLUTION

this is similar to PQ'+PQ

Take A as common we get an expression P(Q+Q')

According to boolean algebra rule Q+Q'=1 hence our expression reduces to P.

In our expression P=(A+B+C), Q=(D+E)

(A+B+C)(D+E)'+(A+B+C)(D+E)

=(A+B+C)((D+E)')+(D+E))

=(A+B+C)(1)

=A+B+C

OPTION A

PLEASE UPVOTE THE SOLUTION